OpenStudy (anonymous):

What is the solution to the system of equations? y = 1.5x – 4 y = –x

OpenStudy (anonymous):

I could definitely be wrong, but from what I just did, I got -3.5. I added 4 to both sides and then subtracted 1.5 from both sides which left me with this... 3.5y = x and b/c y=-x then I changed the final answer to -3.5 I could be wrong, but I tried. Wish ya luck!

OpenStudy (anonymous):

@ratlady16

jhonyy9 (jhonyy9):

so because there is y=-x using this the first equation will be -x=1,5x -4 4=1,5x +x 4=2,5x x = 4/2,5 = 1,6 x=1,6 but you know from second equation that y=-x so this mean that y = - 1,6 hope you understand it

OpenStudy (corpsss):

Make the two equation in the form of x+y=c. Let c be the whole number. Equating the given equation. 1.5x-y=4 x+y=0 Then, using the subtraction and addition of linear equation. This will give you 1.5x=4 x=0 add the two 2.5x=4 (2.5x)/2.5=4/2.5 x= 1.6 Take 1.5x-y=4 -y=4-1.5x substitue x -y=4-1.5(1.6) -y=4-2.4 -y=1.6 divide both side by -1 y=1.6 the solution set are (1.6,1.6) Hope you understand :)

OpenStudy (anonymous):